# The Open Journal of Astrophysics

### v1

@@ 2.3 Relating properties of galaxies to distance
Figure 1 shows the distribution of intrinsic galaxy properties as a function of normalised cluster-centric distance $r/R_{200,\mathrm{mean}}$. Each panel is also overlaid with the median trend line shaded with p1684p16-84 percentiles. The galaxy population regime is described as the following:
Figure 1: Spatial distributions of all fifteen intrinsic galaxy properties. The colour scale indicates number density. Each panel denotes a property, as labelled on the y-axis with the corresponding units if applicable. The x-axis denotes the distance-to-closest-cluster normalised by the $R_{200,\rm mean}$ of the host. Red line denotes the median shaded by 16th-84th percentiles. The plot shows that galaxy properties correlate with galaxy-to-closest-cluster distance.
@@
    Field‐dominated population ($R\gtrsim 5\,R_{200,\mathrm{mean}}$). At larger radii, the distribution coalesces into a more diffused region characterised by high gas‐fractions, elevated sSFR, bluer broad‐band colours, and lower stellar‐to‐halo mass ratios. This population is indicative of isolated, late‐type systems largely unaffected by cluster‐scale environmental mechanisms (e.g. Sulentic et al., 2006; Boselli & Gavazzi, 2006).
The overlaid median curves further quantifies the gradual modulation of galaxy properties with cluster‐centric distance. For instance:
@@
  • Broad‐band colour and metallicity medians shift monotonically, becoming progressively bluer at larger radii. Stellar metallicity becomes less prominent at larger radii, where gas metallicity dominates.
@@ 2.4 Label selection
The extent of a dark matter halo is often defined using a spherical overdensity criterion. $R_{200,\mathrm{crit}}$ is the radius enclosing 200 times the critical density of the Universe ($R_{200\mathrm{c}}$), and $R_{200,\mathrm{mean}}$ the mean matter density ($R_{200,\mathrm{m}}$), where $\rho_{\mathrm{crit}}(z)=3H^{2}(z)/(8\pi G)$ and $\rho_{\mathrm{m}}(z)=\Omega_{\mathrm{m}}(z)\,\rho_{\mathrm{crit}}(z)$. These definitions yield systematically different halo radii, particularly at low redshift (Bryan & Norman, 1998; More et al., 2011; Kravtsov & Borgani, 2012). On the other hand, the splashback radius ($R_{\rm sp}$) provides a dynamically motivated definition of the halo boundary. It corresponds to the radius at which recently accreted material reaches the apocenter of its first orbit, producing a sharp steepening in the spherically averaged density profile (Diemer & Kravtsov, 2014; Adhikari, Dalal & Chamberlain, 2014; More, Diemer & Kravtsov, 2015). Physically, $R_{\rm sp}$ delineates the transition between multi-stream, virialised material in the halo interior and infalling matter on first passage. There are two standard ways to compute $R_{sp}$. The first is the steepest-slope based method introduced in More, Diemer & Kravtsov (2015), which uses a spherically averaged halo density profile and takes the logarithmic slope. The other is SPARTA, outlined in Diemer (2017), which is a particle tracking algorithm that records the first apocentres after infall of all particle orbits and takes the central value to be $R_{sp}$.
@@
In addition to the distance-based label variations, we also incorporate a labelling scheme based on the SubLink merger tree algorithm (Rodriguez-Gomez et al., 2015). We adopt this second method because it is how the simulation “labels” galaxies. A galaxy is considered cluster (label 0) if it has ever been associated with a host group with $M_{200,\mathrm{mean}}>10^{13}\ {\rm M}_{\odot}$ along the main progenitor branch in the merger tree. Otherwise it is a field galaxy (label 1).
@@ 2.5 Data pre-processing: Feature preparation and train-validation-test split
Notes. (1) Mass ratio is the stellar mass $M_{\ast}$ normalised by the total mass of the subhalo $M_{\rm{tot}}$. (2) Gas fraction is defined using the ratio between the mass of gas $M_{\rm gas}$ and the mass of baryons $M_{\rm baryons}$. (3) Specific star formation rate is the SFR in the gas cells divided by stellar mass $M_{\ast}$. (4) Colour is the grg-r magnitude. (5) Gas metallicity is the mass-weighted average metallicity of gas cells $Z_{\rm gas}/Z_{\odot}$. (6) Spin magnitude is the l2l-2 norm of the subhalo’s total spin in the $x$, $y$, and $z$-axis. (7) Star metallicity is the mass-weighted average metallicity of star particles $Z_{\ast}/Z_{\odot}$. (8) The star half-mass radius $R_{\rm\star,1/2}$ is the comoving radius containing half of the subhalo’s stellar mass. (9) Dark-matter mass $M_{\rm DM}$ is defined by the total mass of dark matter matter particles bound to the subhalo. (10) Dark-matter half-mass radius $R_{\rm DM,1/2}$ is the comoving radius containing half of the subhalo’s dark matter mass. (11) Velocity max radius $R_{v_{\rm max}}$ defines the location where $v_{\rm max}$ is achieved. (12) The gas half-mass radius $R_{\rm gas,1/2}$ is comoving radius containing half of the subhalo’s gas mass. (13) Velocity magnitude is $|v|$. (14) Velocity dispersion $\sigma_{v}$ is the one-dimensional velocity dispersion of all particles and cells associated with the subhalo. (15) The max velocity $v_{\rm max}$ is the maximum value of the spherically-averaged rotation curve. Note that sSFR is taken into account as both gas and stellar categories. $\log_{10}$ denotes base-10 logarithm. For arcsinh transforms, parameters $(a,b)$ were chosen to symmetrise each feature’s distribution using the method discuss in Section 2.5.
@@ 2.6 Model construction
Table 2: MLP hyperparameters and detailed layer‐by‐layer parameter counts.
@@ 2.8 Finding the transition region
where $z_{g}$ is the raw logit and $\log_{10}R_{g}$ the logged galaxy-to-cluster distance (normalised by $R_{200,\rm mean}$). Clusters are further grouped into four logarithmic mass bins. These mass ranges are denoted by
$10^{13.0}\leq M_{200\mathrm{m}}/M_{\odot}<10^{13.5},\qquad 10^{13.5}\leq M_{200\mathrm{m}}/M_{\odot}<10^{14.0},$
$10^{14.0}\leq M_{200\mathrm{m}}/M_{\odot}<10^{14.5},\qquad 10^{14.5}\leq M_{200\mathrm{m}}/M_{\odot}<10^{15.0}.$
We aim to assign each cluster mass bin a single zero–crossing radius, which we compute in parallel by two complementary schemes: taking the cluster-to-cluster median of individual cluster zero points and finding zero points on stacked profiles, written as $r^{\rm med}$ and $r^{\rm stack}$ respectively. All zero-point computations were performed separately within each mass range to reveal mass‐dependence. We describe these schemes in more detail below.
@@
Figure 3: Per-mass bin median probability profile of the fiducial model using distance-based labels ($0.5/5.0$). The shaded region represents the 16-84 percentile scattering around the median. The y-axis denotes the raw logits predicted by the model for each galaxy sample. The $x$-axis denotes distance normalised by $R_{200,\rm mean}$. Logits are first collected to make the per-cluster profile based on the host cluster of each galaxy, which is then stacked into mass bins for the final profile. This clearly demonstrates the probabilistic relationship between the normalised galaxy-to-cluster distance and galaxy membership (cluster or field).
@@ 3.1 Convergence test: Detailed discussion
A convergence test is performed by varying the upper distance ratio threshold of field galaxy labels, using the same model construction in Section 2.6. The upper threshold is incremented by 0.1, creating 41 labelled sets in total. We apply the convergence test on all feature combinations, where the model is shown to converge regardless of the features we choose.
@@
We now evaluate and discuss the outcome of the convergence test. This verifies our model’s predicted cluster transition radius are robust to the different choices of distance‐based labelling. The number of labelled samples also varies based on the labels. For each subset of trained models, we fix the network architecture, optimiser settings (learning rate schedule, weight decay, batch size), and random seed governing both weight initialization and data shuffling. We adopt the same train/validation/test split for all experiments, with early‐stopping triggered by the plateau in validation loss to avoid overfitting. Each training run yields a predicted transition radius, defined as the radial location where the median cluster-to-cluster zero‐crossing occurs for every mass bin, together with the 16th and 84th percentiles to represent the inherent scatter. The outcome demonstrates that in each mass bin, the prediction region (median and p-16 84 errors) does not increase as we increase the upper threshold. This shows that the model is successful at determining a spatial region of separation among the two memberships, where the difference between the predicted transition regions decreases as the upper threshold goes beyond $4$. This is also consistent regardless of the feature set the model is trained on.
@@
Figure 4 shows the median transition radii for a single mass bin $10^{14.0}\leq M_{200,\mathrm{mean}}/M_{\odot}<10^{14.5}$, plotted as a function of the incremental upper distance threshold for field galaxies. Each column denotes a different radii definition used for labelling ($r/R_{sp}$, $r/R_{200,\rm mean}$, and $r/R_{200,\rm crit}$ respectively). The same trend is observed across the other mass bins. The top row represents models trained using the six primary properties, and the bottom uses all fifteen properties. Even when the threshold increases more than five times, the shifts in the inferred radius never exceed the width of the 16–84% bands—typically $\Delta(r/R_{200m})\lesssim 0.5$. This demonstrates that our model’s prediction of the transition region is not an artifact of a particular distance-based label threshold: once a split of cluster-like and field-like galaxies are established through the labelled sets, the transition region becomes a persistent identity. Although models labelled using all three radii definitions exhibit convergence behaviour, the $R_{sp}$-based models have the most stable trend, while successfully demonstrating the expected intrinsic scattering. Finally, the same convergence behaviour is observed across different feature subsets (e.g. six‐property, all‐property, gas‐only, stellar‐only, and dynamics only).
@@ 3.2 Spatial extent of the cluster environment’s influence
Distances are expressed as $r/R_{200,\rm mean}$, with $R_{200,\rm mean}$ defined for the host cluster of each galaxy. All host clusters are distributed across four mass bins. The results are produced by the fiducial distance based model denoted by $0.5/5.0$ and the SubLink based model. For a given model and mass bin, we compute (1) the halo-level $z(r)$ profiles, binned in $\log r$, and (2) the median and 16–84th percentile envelopes across cluster mass bins. Figure 3 displays the median $z(r)$ curve of model $0.5/5.0$. This probability profile has the following morphology: an inner region where $z<0$, followed by a rapid rise between $\sim 0.5$ and $\sim 1.5\,R_{200,\mathrm{mean}}$ where the distribution passes $z=0$, and an outer region beyond $\sim 2\,R_{200,\mathrm{mean}}$ where $z$ asymptotically approaches positive values. The distributions of lower-mass hosts tend to show slightly higher $z$ at fixed radius inside $R_{200,\mathrm{mean}}$. Overall the mass bin distributions exhibit uniformity across different distance based models.
@@ 3.3 Intrinsic scatter
The bright yellow curve delineates the normalised spatial distribution of confident field galaxies. These have very positive logits in $5.0,\infty]$, and their distance distribution is nearly zero inside $1\,R_{200,\mathrm{mean}}$, then gradually rises beyond.
@@
It is important to note that the dark purple line and the bright yellow line intersect slightly beyond $R_{200,\rm mean}$. This indicates that this region contains a fraction of both confident cluster and field members, further suggesting that the galaxy membership transition likely takes place here. This suggests that the set of galaxy properties we feed into the network (stellar-to-subhalo mass ratio, gas fraction, colour, and metallicities…) exhibit gradual changes across the $R_{200,\rm mean}$. Each physical category of properties also experience change at different locations and under various time scales, due to the complexity in quenching mechanisms galaxies may experience. This is explained in detail in Section 3.6.
@@
Figure 6: Radial distribution of galaxy logit output ranges from the fiducial model ($0.5/5.0$ with six properties). The curves show the normalised contribution of galaxies within a sequence of logit bins as a function of $\log(r/R_{200,\mathrm{mean}})$. The colour bar denotes the boundary of bins. There are $\infty,-5.0]$ on the far left end (purple) and $5.0,\infty]$ on the far right end (yellow). In between there are 21 logit bins with inner step $\approx 0.4$. The logit bin [0.23,0.23][-0.23,0.23] encloses the zero point, which is demarcated in the middle. Therefore, the dark purple curve indicates confidently cluster-like samples ($\mathrm{logit}\in[-\infty,-5.0]$), and yellow represents field-like samples ($\mathrm{logit}\in[5.00,\infty]$). In between, the logit bins show intrinsic scattering of galaxy samples, where [0.23,0.23][-0.23,0.23] contains the most ambiguous sample. This shows that there is intrinsic scattering of ambiguous galaxy samples across all radial bins, indicating that galaxies are impacted by the cluster environment through a gradual process.
@@ 3.4 Mass dependence of the transition region
According existing theory, dark‐matter haloes exhibit a well‐established mass-concentration relation of the form $c(M)\propto M^{-\alpha}$ with $\alpha\approx 0.1$ (e.g., Bullock et al., 2001; Dolag, K. et al., 2004; Dutton & Macciò, 2014, etc.). As haloes of different masses collapse at different epochs in hierarchical structure formation, lower-mass haloes typically form earlier, when the mean cosmic density is higher. This leads to more compact central regions, indicating higher concentrations. Their high-mass counterparts, however, tend to assemble later, producing shallower inner profiles. In the NFW formalism, this means that massive haloes have larger scale radius $r_{s}$, which implies that their density profiles decline more gradually with radius compared to low-mass haloes Navarro, Frenk & White (1996). Consequently, lower‐mass haloes are on average more centrally concentrated than their higher‐mass counterparts.

### v2

@@ 2.3 Relating properties of galaxies to distance
Figure 1 shows the distribution of intrinsic galaxy properties as a function of normalised cluster-centric distance $r/R_{200,\mathrm{mean}}$. Each panel is also overlaid with the median trend line shaded with the 16th – 84th percentiles. The galaxy population regime is described as the following:
Figure 1: Spatial distributions of all fifteen intrinsic galaxy properties. The colour scale indicates number density. Each panel denotes a property, as labelled on the y-axis with the corresponding units if applicable. The x-axis denotes the distance-to-closest-cluster normalised by the $R_{200,\rm mean}$ of the host. Red line denotes the median shaded by 16th-84th percentiles. The plot shows that galaxy properties correlate with galaxy-to-closest-cluster distance. The list of properties is specified in Table 1.
@@
    Field‐dominated population ($R\gtrsim 3\,R_{200,\mathrm{mean}}$). At larger radii, the distribution coalesces into a more diffused region characterised by high gas‐fractions, elevated sSFR, bluer broad‐band colours, and lower stellar‐to‐halo mass ratios. It is important to note that the galaxies between $3\,R_{200,\mathrm{mean}}$ and $5\,R_{200,\mathrm{mean}}$ of their host cluster do exhibit field-like properties, even though the trend-lines stabilize beyond $5\,R_{200,\mathrm{mean}}$. These populations are indicative of isolated, late‐type systems largely unaffected by cluster‐scale environmental mechanisms (e.g. Sulentic et al., 2006; Boselli & Gavazzi, 2006).
The overlaid median curves further quantify the gradual modulation of galaxy properties with cluster‐centric distance. For instance:
@@
  • Broad‐band colour and metallicity medians shift monotonically, becoming progressively bluer at larger radii of the cluster. Stellar metallicity becomes less prominent at larger cluster radii, where gas metallicity dominates.
@@ 2.4 Label selection
The extent of a dark matter halo is often defined using a spherical overdensity criterion. $R_{200,\mathrm{crit}}$ is the radius enclosing 200 times the critical density of the Universe ($R_{200\mathrm{c}}$), and $R_{200,\mathrm{mean}}$ the mean matter density ($R_{200,\mathrm{m}}$), where $\rho_{\mathrm{crit}}(z)=3H^{2}(z)/(8\pi G)$ and $\rho_{\mathrm{m}}(z)=\Omega_{\mathrm{m}}(z)\,\rho_{\mathrm{crit}}(z)$. These definitions yield systematically different halo radii, particularly at low redshift (Bryan & Norman, 1998; More et al., 2011; Kravtsov & Borgani, 2012). On the other hand, there are two standard ways to compute $R_{sp}$. The first is the steepest-slope based method introduced in More, Diemer & Kravtsov (2015), which uses a spherically averaged halo density profile and takes the logarithmic slope. The other is SPARTA, outlined in Diemer (2017), which is a particle tracking algorithm that records the first apocentres after infall of all particle orbits and takes the central value to be $R_{sp}$.
@@
In addition to the distance-based label variations, we also incorporate a labelling scheme based on the SubLink merger tree algorithm (Rodriguez-Gomez et al., 2015). We adopt this second method because it is how the simulation “labels” galaxies. A galaxy is considered cluster, label 0, if it has ever been associated with a host group with $M_{200,\mathrm{mean}}>10^{13}\ {\rm M}_{\odot}$ along the main progenitor branch in the merger tree. Otherwise it is a field galaxy (label 1).
@@ 2.5 Data pre-processing: Feature preparation and train-validation-test split
Notes. (1) Mass ratio is the stellar mass $M_{\ast}$ normalised by the total mass of the subhalo $M_{\rm{tot}}$. (2) Gas fraction is defined using the ratio between the mass of gas $M_{\rm gas}$ and the mass of baryons $M_{\rm baryons}$. (3) Specific star formation rate is the SFR in the gas cells divided by stellar mass $M_{\ast}$. (4) Colour is the grg-r magnitude. (5) Gas metallicity is the mass-weighted average metallicity of gas cells $Z_{\rm gas}/Z_{\odot}$. (6) Spin magnitude is the l2l-2 norm of the subhalo’s total spin in the $x$, $y$, and $z$-axis. (7) Star metallicity is the mass-weighted average metallicity of star particles $Z_{\ast}/Z_{\odot}$. (8) The star half-mass radius $R_{\rm\star,1/2}$ is the comoving radius containing half of the subhalo’s stellar mass. (9) Dark-matter mass $M_{\rm DM}$ is defined by the subhalo mass type 1, summing the mass of all dark matter particles bound to the subhalo. (10) Dark-matter half-mass radius $R_{\rm DM,1/2}$ is the comoving radius containing half of the subhalo’s dark matter mass. (11) Velocity max radius $R_{v_{\rm max}}$ defines the location where $v_{\rm max}$ is achieved. (12) The gas half-mass radius $R_{\rm gas,1/2}$ is comoving radius containing half of the subhalo’s gas mass. (13) Velocity magnitude is $|v|$. (14) Velocity dispersion $\sigma_{v}$ is the one-dimensional velocity dispersion of all particles and cells associated with the subhalo. (15) The max velocity $v_{\rm max}$ is the maximum value of the spherically-averaged rotation curve. Note that sSFR is taken into account as both gas and stellar categories. $\log_{10}$ denotes base-10 logarithm. For arcsinh transforms, parameters $(a,b)$ were chosen to symmetrise each feature’s distribution using the method discuss in Section 2.5.
@@ 2.6 Model construction
Table 2: MLP hyperparameters and detailed layer-by-layer parameter counts.
@@ 2.8 Finding the transition region
where $z_{g}$ is the raw logit and $\boldsymbol{\ell}_{h}$ the logged galaxy-to-cluster distance (normalised by $R_{200,\rm mean}$). As previously stated, clusters are grouped into four logarithmic mass bins. We aim to assign each cluster mass bin a single zero–crossing radius, which we compute in parallel by two complementary schemes: taking the cluster-to-cluster median of individual cluster zero points and finding zero points on stacked profiles, written as $r^{\rm med}$ and $r^{\rm stack}$ respectively. All zero-point computations were performed separately within each mass range to reveal mass‐dependence. We describe these schemes in more detail below.
@@
Figure 3: Per-mass bin median probability profile of the fiducial model using distance-based labels ($0.5/5.0$). The shaded region represents the 16-84 percentile scattering around the median. The y-axis is the raw logits (z) predicted by the model for each galaxy sample. Retaining the raw logits for the y-axis allows the flattening at two ends to be demonstrated more clearly. The $x$-axis denotes distance normalised by $R_{200,\rm mean}$. Logits are first collected to make the per-cluster profile based on the host cluster of each galaxy, which is then stacked into mass bins for the final profile. This clearly demonstrates the probabilistic relationship between the normalised galaxy-to-cluster distance and galaxy membership (cluster or field).
@@ 3.1 Convergence test: Detailed discussion
A convergence test is performed by varying the upper distance ratio threshold of field galaxy labels, using the same model construction as in Section 2.6. The upper threshold is incremented by 0.1, creating 41 labelled sets in total. We apply the convergence test on all feature combinations, where the model is shown to converge regardless of the features we choose.
@@
We now evaluate and discuss the outcome of the convergence test. This verifies our model’s predicted cluster transition radius are robust to the different choices of distance‐based labelling. The number of labelled samples also varies based on the labels. For each subset of trained models, we fix the network architecture, optimiser settings (learning rate schedule, weight decay, batch size), and random seed governing both weight initialization and data shuffling. We adopt the same train/validation/test split for all experiments, with early‐stopping triggered by the plateau in validation loss to avoid overfitting. Each training run yields a predicted transition radius, defined as the radial location where the median cluster-to-cluster zero‐crossing occurs for every mass bin, together with the 16th and 84th percentiles to represent the inherent scatter. The outcome demonstrates that in each mass bin, the prediction region (median and 16th and 84th percentile errors) does not increase as we increase the upper threshold. This shows that the model is successful at determining a spatial region of separation among the two memberships, where the difference between the predicted transition regions decreases as the upper threshold goes beyond $4$. This is also consistent regardless of the feature set the model is trained on.
@@
Figure 4 shows the median transition radii for a single mass bin $10^{14.0}\leq M_{200,\mathrm{mean}}/M_{\odot}<10^{14.5}$, plotted as a function of the incremental upper distance threshold for field galaxies. Each column denotes a different radii definition used for labelling ($r/R_{sp}$, $r/R_{200,\rm mean}$, and $r/R_{200,\rm crit}$ respectively). The same trend is observed across the other mass bins. The top row represents models trained using the six primary properties, and the bottom uses all fifteen properties. Even when the threshold increases more than five times, the shifts in the inferred radius never exceed the width of the 16–84% bands—typically $\Delta(r/R_{200m})\lesssim 0.5$. This demonstrates that our model’s prediction of the transition region is not an artifact of a particular distance-based label threshold: once a split of cluster-like and field-like galaxies are established through the labelled sets, the transition region becomes a persistent identity. Although models labelled using all three radii definitions exhibit convergence behaviour, the $R_{sp}$-based models have the most stable trend, while successfully demonstrating the expected intrinsic scattering. Finally, the same convergence behaviour is observed across different feature subsets (e.g. six‐property, all‐property, gas‐only, stellar‐only, and dynamics only). These feature combinations are added to test whether different property combinations exhibit changes at different locations.
@@ 3.2 Spatial extent of the cluster environment’s influence
Distances are expressed as $r/R_{200,\rm mean}$, with $R_{200,\rm mean}$ defined for the host cluster of each galaxy. All host clusters are distributed across four mass bins. The results are produced by the fiducial distance based model denoted by $0.5/5.0$ and the SubLink based model. For a given model and mass bin, we compute (1) the halo-level $z(r)$ profiles, binned in $\log r$, and (2) the median and 16-84th percentile envelopes across cluster mass bins. Figure 3 displays the median $z(r)$ curve of model $0.5/5.0$. This probability profile has the following morphology: an inner region where $z<0$, followed by a rapid rise between $\sim 0.5$ and $\sim 1.5\,R_{200,\mathrm{mean}}$ where the distribution passes $z=0$, and an outer region beyond $\sim 2\,R_{200,\mathrm{mean}}$ where $z$ asymptotically approaches positive values. The distributions of lower-mass hosts tend to show slightly higher $z$ at fixed radius inside $R_{200,\mathrm{mean}}$. Overall the mass bin distributions exhibit uniformity across different distance based models.
@@ 3.3 Intrinsic scatter
The light green curve delineates the normalised spatial distribution of confident field galaxies. These have very positive logits in $5.0,\infty]$, and their distance distribution is nearly zero inside $1\,R_{200,\mathrm{mean}}$, then gradually rises beyond.
@@
It is important to note that the dark purple line and the light green line intersect slightly beyond $R_{200,\rm mean}$. This indicates that this region contains a fraction of both confident cluster and field members, further suggesting that the galaxy membership transition likely takes place here. This suggests that the set of galaxy properties we feed into the network (stellar-to-subhalo mass ratio, gas fraction, colour, and metallicities…) exhibit gradual changes across the $R_{200,\rm mean}$. Each physical category of properties also experience change at different locations and under various time scales, due to the complexity in quenching mechanisms galaxies may experience. This is explained in detail in Section 3.6.
@@
Figure 6: Radial distribution of galaxy logit output ranges from the fiducial model ($0.5/5.0$ with six properties). The curves show the normalised contribution of galaxies within a sequence of logit bins as a function of $\log(r/R_{200,\mathrm{mean}})$. The colour bar denotes the boundary of bins. There are $\infty,-5.0]$ on the far left end (purple) and $5.0,\infty]$ on the far right end (light green). In between there are 21 logit bins with inner step $\approx 0.4$. The logit bin [0.23,0.23][-0.23,0.23] encloses the zero point, which is demarcated in the middle. Therefore, the dark purple curve indicates confidently cluster-like samples ($\mathrm{logit}\in[-\infty,-5.0]$), and light green represents field-like samples ($\mathrm{logit}\in[5.00,\infty]$). In between, the logit bins show intrinsic scattering of galaxy samples, where [0.23,0.23][-0.23,0.23] contains the most ambiguous sample. This shows that there is intrinsic scattering of ambiguous galaxy samples across all radial bins, indicating that galaxies are impacted by the cluster environment through a gradual process.
@@ 3.4 Mass dependence of the transition region
According to existing theory, dark‐matter haloes exhibit a well‐established mass-concentration relation of the form $c(M)\propto M^{-\alpha}$ with $\alpha\approx 0.1$ (e.g., Bullock et al., 2001; Dolag, K. et al., 2004; Dutton & Macciò, 2014, etc.). As haloes of different masses collapse at different epochs in hierarchical structure formation, lower-mass haloes typically form earlier, when the mean cosmic density is higher. This leads to more compact central regions, indicating higher concentrations. Their high-mass counterparts, however, tend to assemble later, producing shallower inner profiles. In the NFW formalism, this means that massive haloes have larger scale radius $r_{s}$, which implies that their density profiles decline more gradually with radius compared to low-mass haloes (Navarro, Frenk & White, 1996). Consequently, lower‐mass haloes are on average more centrally concentrated than their higher‐mass counterparts.